# Efficient Training-Free Online Routing for High-Volume Multi-LLM Serving

### v2

@@ Appendix E Extended Related Work
LLM Routing. Existing LLM routing research primarily falls into two paradigms. The first focuses on improving response quality while managing cost, typically through ensembling outputs from different LLMs [26, 55] or using cascading strategies that query LLMs sequentially based on their capabilities [5, 2, 62, 33]. For instance, LLM-Blender [26] proposes an ensembling framework that combines outputs from multiple open-source LLMs and selects the optimal response.  [55]using outputs Wang et al. [55] propose to fuse outputs of expert models that capture complementary aspects of the data distribution, to generate the final answer. Frugal-GPT [5] adopts a sequential querying strategy, invoking LLMs in order of increasing capability until a satisfactory response is obtained. AutoMix [2] first uses a smaller model to self-verify the quality of the response, based on which it dynamically selects a suitably sized model for handling the query. Although these approaches can improve performance, they incur high latency and computational cost due to multiple model invocations per query.
@@
The second paradigm leverages historical or auxiliary datasets to estimate the performance and cost of each query and select the optimal LLM accordingly [43, 50, 23, 13, 37, 27, 48, 52, 15, 20, 56]. In this line of work, several approaches train model-based predictors using these labeled historical datasets. For example, RouterLLM [43] trains a BERT-base or causal LLM on the historical data annotated with human preferences to enhance routing accuracy. CARROT [50] trains Roberta-base models to estimate both performance and cost for each query, enabling selection of the optimal LLM under any user-defined trade-off between quality and cost. HybridLLM [13] adopts the synthetic preference labels from the MixInstruct dataset [26] based on BARTScore [61], and trains a single BERT-based router for query routing. Zooter [37] distills reward signals from training queries and applies tag-based label enhancement to train a routing function that facilitates expert selection. TensorOpera [52] introduces a soft label-based strategy based on the BERT similarity scores to train the BERT-based predictor for query routing. GraphRouter [15] constructs a heterogeneous graph to capture the contextual relationship between the query requirements and the LLM capabilities for informed routing decisions. While these training-based approaches are effective for certain LLM deployment settings, they introduce nontrivial training overhead. Furthermore, adapting them to varying LLM deployment configurations typically requires retraining, which makes them unsuitable for resource-constrained online routing with diverse LLM deployment configurations.
@@
Some methods avoid training by using approaches such as KNN [23] or similarity-weighted (SW) ranking [43] to estimate the performance and cost based on the historical data. While these approaches avoid model training overhead, they still incur substantial computational overhead and latency, as they rely on brute-force comparisons against the entire historical dataset to retrieve similar examples or compute similarity scores without any optimization. Specifically, they operate with a high search complexity of $O(N)$, where $N$ is the size of the dataset. As a result, these methods are difficult to scale and impractical for high-volume, budget-constrained online routing scenarios.
@@
Recent efforts [47, 42, 44] have explored formulating LLM routing as MILP. For instance, Sakota et al. [47] introduce two MILP-based strategies, targeting performance-oriented and cost-oriented, respectively. However, these methods face significant challenges in high-volume online routing settings, where queries arrive sequentially, rather than simultaneously, making the offline optimal infeasible to compute.
@@
Online Matching. This problem has been extensively studied for decades, with much of the literature focusing on important special cases of the general online submodular welfare maximization problem [29, 32]. These include the classical unweighted online bipartite matching problem [28, 30, 31], its extension to vertex-weighted matching [4, 1], the display ads allocation problem [14, 6], and the AdWords problem [39, 11, 4].

### v3

@@ Appendix E Extended Related Work
LLM Routing. Existing LLM routing research primarily falls into two paradigms. The first focuses on improving response quality while managing cost, typically through ensembling outputs from different LLMs [jiang2023llm, wang2023fusing] or using cascading strategies that query LLMs sequentially based on their capabilities [chen2023frugalgpt, aggarwal2024automix, yue2023large, lee2023orchestrallm]. For instance, LLM-Blender [jiang2023llm] proposes an ensembling framework that combines outputs from multiple open-source LLMs and selects the optimal response.  [wang2023fusing]using outputs Wang et al. [wang2023fusing] propose to fuse outputs of expert models that capture complementary aspects of the data distribution, to generate the final answer. Frugal-GPT [chen2023frugalgpt] adopts a sequential querying strategy, invoking LLMs in order of increasing capability until a satisfactory response is obtained. AutoMix [aggarwal2024automix] first uses a smaller model to self-verify the quality of the response, based on which it dynamically selects a suitably sized model for handling the query. Although these approaches can improve performance, they incur high latency and computational cost due to multiple model invocations per query.
@@
The second paradigm leverages historical or auxiliary datasets to estimate the performance and cost of each query and select the optimal LLM accordingly [ong2025routellmlearningroutellms, somerstep2025carrot, hu2024routerbench, ding2024hybrid, lu2023routing, jitkrittum2025universal, shnitzer2023large, stripelis2024tensoropera, feng2024graphrouter, hari2023tryage, wang2025mixllm]. In this line of work, several approaches train model-based predictors using these labeled historical datasets. For example, RouterLLM [ong2025routellmlearningroutellms] trains a BERT-base or causal LLM on the historical data annotated with human preferences to enhance routing accuracy. CARROT [somerstep2025carrot] trains Roberta-base models to estimate both performance and cost for each query, enabling selection of the optimal LLM under any user-defined trade-off between quality and cost. HybridLLM [ding2024hybrid] adopts the synthetic preference labels from the MixInstruct dataset [jiang2023llm] based on BARTScore [yuan2021bartscore], and trains a single BERT-based router for query routing. Zooter [lu2023routing] distills reward signals from training queries and applies tag-based label enhancement to train a routing function that facilitates expert selection. TensorOpera [stripelis2024tensoropera] introduces a soft label-based strategy based on the BERT similarity scores to train the BERT-based predictor for query routing. GraphRouter [feng2024graphrouter] constructs a heterogeneous graph to capture the contextual relationship between the query requirements and the LLM capabilities for informed routing decisions. While these training-based approaches are effective for certain LLM deployment settings, they introduce nontrivial training overhead. Furthermore, adapting them to varying LLM deployment configurations typically requires retraining, which makes them unsuitable for resource-constrained online routing with diverse LLM deployment configurations.
@@
Some methods avoid training by using approaches such as KNN [hu2024routerbench] or similarity-weighted (SW) ranking [ong2025routellmlearningroutellms] to estimate the performance and cost based on the historical data. While these approaches avoid model training overhead, they still incur substantial computational overhead and latency, as they rely on brute-force comparisons against the entire historical dataset to retrieve similar examples or compute similarity scores without any optimization. Specifically, they operate with a high search complexity of $O(N)$, where $N$ is the size of the dataset. As a result, these methods are difficult to scale and impractical for high-volume, budget-constrained online routing scenarios.
@@
Recent efforts [vsakota2024fly, mohammadshahi2024routoo, rama2025cerebrum] have explored formulating LLM routing as MILP. For instance, Sakota et al. [vsakota2024fly] introduce two MILP-based strategies, targeting performance-oriented and cost-oriented, respectively. However, these methods face significant challenges in high-volume online routing settings, where queries arrive sequentially, rather than simultaneously, making the offline optimal infeasible to compute.
@@
Online Matching. This problem has been extensively studied for decades, with much of the literature focusing on important special cases of the general online submodular welfare maximization problem [kapralov2013online, korula2015online]. These include the classical unweighted online bipartite matching problem [kalyanasundaram2000optimal, karp1990optimal, kesselheim2013optimal], its extension to vertex-weighted matching [buchbinder2007online, aggarwal2011online], the display ads allocation problem [feldman2009online, chen2011real], and the AdWords problem [mehta2007adwords, devanur2009adwords, buchbinder2007online].